OpenStudy (anonymous):

Caitlin earns $7.00 an hour doing yard work for her neighbor. She worked 3 hours Friday, 4 hours Saturday, and more hours Sunday. She earned $77.00 for the weekend. How many hours did she do yard work on Sunday? A. 4 B. 5 C. 6 D. 7

OpenStudy (lucaz):

$7/h 7(3)+7(4)+7(sunday)=77 7(s)=77-49 (s)=28/7 hours sunday=4